A researcher is interested in determining the average number of years employees of a company stay with the company. If past information shows a standard deviation of 6 months, what sample size should be taken so that, at 95% confidence, the margin of error will be 4 months or less. Assume that the data is normally distributed.


1
Expert's answer
2022-03-11T11:43:17-0500

"ME=z_{0.025}\frac{\sigma}{\sqrt{n}}."

"n=(\frac{z_{0.025}\sigma}{ME})^2=(\frac{1.96*6}{4})^2=9."